OpenStudy (anonymous):

Lucy’s family purchased one sweat shirt and four tee shirts for a total of $110. Mark’s family picked up three sweatshirts and four tee shirts for a total of $200. Write algebra equations for each purchase in the same way you did in problem one. Use your equations to find the cost of a tee shirt and a sweat shirt.

OpenStudy (hoblos):

cost of sweat shirt =x cost of a tee shirt = y x+4y = 110 3x+4y = 200

OpenStudy (hoblos):

do you know how to solve it now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x+4y do i solve this equation?

OpenStudy (hoblos):

well you have to solve the whole system... try subtracting the two equations.. what will you get?

OpenStudy (anonymous):

i got 110-4y

OpenStudy (hoblos):

x+4y = 110 --------- equation (1) 3x+4y = 200 --------- equation (2) if you subtract you will get